The healthcare provider writes an order for levothyroxine sodium 50 micrograms PO stat. Each tablet contains 0.05 mg. How many tablets would you administer?

A.1 tablet

B. 2 tablets

C.1/2 tablet (half) a tablet

D.3 tablets
Answer #1

Answer: A. 1 tablet

Explanation: Since each tablet contains=0.05 mg=0.05*10-3 g=5*10-5 grams. [1 mg=10-3 g].

Needed quantity=50 micrograms=50*10-6 g =5*10-5 grams. [1 microgram=10-6 g].

So, Number of tablets=(Needed quantity)/(quantity contained in 1 tablet)=(5*10-5 g)/(5*10-5 g)=1 tablet.